OpenStudy (anonymous):

(x^2)*(e^x)-(5)(x)(e^x)=0

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

x^2e^x-5xe^x=0 ... Start with the given equation. xe^x(x-5)=0 ... Factor out the GCF x=0 or e^x=0 or x-5=0 ... Use the zero product property x=0 or x=5 ... Solve for x Note: there are no solutions to e^x=0, so we ignore this term. So the solutions are x=0 or x=5